### v3

# Faster and Better LLMs via Latency-Aware Test-Time Scaling

Abstract
Test-Time Scaling (TTS) has proven effective in improving the performance of Large Language Models (LLMs) during inference. However, existing research has overlooked the efficiency of TTS from a latency-sensitive perspective. Through a latency-aware evaluation of representative TTS methods, we demonstrate that a compute-optimal TTS does not always result in the lowest latency in scenarios where latency is critical. To address this gap and achieve latency-optimal TTS, we propose two key approaches by optimizing the concurrency configurations: (1) branch-wise parallelism, which leverages multiple concurrent inference branches, and (2) sequence-wise parallelism, enabled by speculative decoding. By integrating these two approaches and allocating computational resources properly to each, our latency-optimal TTS enables a 32B model to reach 82.3% accuracy on MATH-500 within 1 minute and a smaller 3B model to achieve 72.4% within 10 seconds. Our work emphasizes the importance of latency-aware TTS and demonstrates its ability to deliver both speed and accuracy in latency-sensitive scenarios.
Faster and Better LLMs via Latency-Aware Test-Time Scaling
Zili Wang1,2thanks: Equal contribution., Tianyu Zhang3∗, Lei Zhu3,†, Haoli Bai3, Lu Hou3, Shiming Xiang1,2,†, Xianzhi Yu3, Wulong Liu3 1 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 2 Institute of Automation, Chinese Academy of Sciences, China 3 Huawei Noah’s Ark Lab wangzili2022@ia.ac.cn, {zhangtianyu59,zhulei168}@huawei.com
Figure 1: “Compute-optimal” does not necessarily translate to “latency-optimal” for test-time scaling. Left: Previous works measure test-time scaling by #token budget, indicating sequential scaling shows superior token efficiency than parallel scaling (majority voting for instance). Right: When considering latency as budget, parallel scaling can be 1.6x faster to achieve the same accuracy than sequential scaling. Experiments are performed using s1.1-32B Muennighoff et al. (2025) on MATH-500 Hendrycks et al. (2021). The red curve corresponds to sequential scaling with sequence length varying in {1024, 2048, 4096, 8192}, via budget forcing. The blue curve represents parallel scaling by majority voting with a fixed sequence length.
   Equal Contributions.   Corresponding Author.
Figure 2: Latency-aware test-time scaling on MATH-500 with different model types, with sequential scaling in red and parallel scaling in blue.
1 Introduction
Test-Time Scaling (TTS) is an effective approach to improve the performance of Large Language Models (LLMs) at the cost of additional inference-time computations Snell et al. (2024); Brown et al. (2024). TTS can be realized by two basic approaches: sequential scaling and parallel scaling. Sequential scaling requires the model to produce an extended reasoning process in a single pass Muennighoff et al. (2025). In contrast, parallel scaling generates multiple solutions in parallel and selects the final answer, usually through majority voting Wang et al. (2022); Liu et al. (2025). Hybrid approaches can be constructed on top of the two basic ones Guan et al. (2025); Wang et al. (2024b).
With the number of generated tokens (#tokens) as budget, many existing studies Snell et al. (2024); Setlur et al. (2025); Yang et al. (2025); Liu et al. (2025); Shi and Jin (2025); Zhang et al. (2024) have examined compute-optimal strategies that enhance average performance gain per token, a metric we refer to as token efficiency. However, in latency-sensitive scenarios where small batch sizes are employed, “compute-optimal” does not necessarily translate to “latency-optimal”. This discrepancy is because the performance achieved within a limited time is determined by token efficiency as well as the throughput (i.e., average number of output tokens per second). As shown in Figure 1, for s1.1-32B Muennighoff et al. (2025) model, although sequential scaling has a better token efficiency, achieving a similar performance with 9$\times$ fewer tokens compared to parallel scaling, it turns out that parallel scaling achieves 1.6$\times$ lower latency. In fact, under small batch sizes, the time of an LLM autoregressive decoding step is dominated by the memory access of the parameters. Therefore, a moderately increased number of parallel branches incurs little additional latency, allowing a much higher throughput almost for free, as shown in Figure 3.
Figure 3: System state, latency, and throughput of s1.1-32B. Left: The roofline model with s1.1-32B. Increasing computational demand shifts execution from memory-bound to compute-bound. Middle: Throughput scales linearly with batch size before saturating at peak FLOPS. Right: The latency per forward pass under varying batch sizes.
We then ask how we can achieve latency-optimal test-time scaling. One lesson from the observation above is that, in addition to optimizing token efficiency, attention must be given to improving the generation concurrency to chase a throughput. There are two approaches to improving concurrency: (1) branch-wise parallelism, which increases the number of parallel branches $B$, and (2) sequence-wise parallelism, which generates multiple successive tokens for a sequence in a single forward pass with speculative decoding Leviathan et al. (2023); Chen et al. (2023). However, current research lacks a thorough analysis of how to allocate computational power to these two resource-competing approaches and to what extent they can be improved.
To bridge the gap, we examine the impact of concurrency configuration for latency-aware test-time scaling in this study. Specifically, we conduct experiments on representative datasets including MATH-500 Hendrycks et al. (2021), AIME24 AoPS (2024), AIME25 AoPS (2025), and GPQA-Diamond Rein et al. (2024) across model sizes from 3B to 32B and under varied concurrency configurations. Revealed by our experiments and analyses, the latency-optimal concurrency depends on the comparative advantage of token efficiency in sequential scaling and parallel scaling. For models with higher token efficiency with parallel scaling (e.g., LLaMa-3.1-8B-Instruct Meta (2024)), one should prioritize branch-wise parallelism. Otherwise (e.g., for QwQ-32B Team (2025)) sequence-wise parallelism takes priority. Besides, under a latency constraint of 1 minute, an optimal concurrency configuration can provide up to 7.3% better performance than baselines where only a single parallelism is applied.
Our contributions can be summarized as follows:
  1. 1.
    We introduce latency-aware test-time scaling, which considers the performance scaling under latency constraints. Unlike existing works that prominently only focus on token efficiency, we discover the necessity of taking systemic throughput into consideration for a better latency-performance trade-off.
  2. 2.
    We provide a unified view for parallel branches and speculative decoding from the perspective of generation concurrency. This allows us to frame latency-optimal test-time scaling as a resource allocation problem.
  3. 3.
    Through extensive experiments, we explore the optimal concurrency configuration for latency-aware test-time scaling. Our experiments reveal that for s1.1-32B, an optimal concurrency configuration can improve the accuracy by 7.3% while reducing latency by 1.7$\times$, reaching 82.3% on MATH-500 in 1 minute.
2 Related Work
Test-time scaling
Scaling the compute at inference time has been proven to be a prominent approach to improve the performance of LLMs. Generally, test-time scaling methods fall into two main categories: sequential and parallel scaling. For the former, sequential scaling methods, represented by OpenAI’s o1 OpenAI (September 2024), DeepSeek-R1 Guo et al. (2025), and Qwen QwQ Team (2025), enforce the model to generate longer solutions with a detailed reasoning chain. Such an ability can be incentivized with supervised training Nye et al. (2021); Lee et al. (2025); Muennighoff et al. (2025) or reinforcement learning Guo et al. (2025). Another main category is parallel scaling. This method first generates multiple response candidates in parallel, then applies a selection criterion to identify the best output. Recent research primarily uses token count as a metric to measure the budget of test-time scaling Snell et al. (2024); Setlur et al. (2025); Yang et al. (2025). Different from previous works, our work points out that compute-optimal does not necessarily translate to latency-optimal. We introduce latency-aware test-time scaling and figure out the concurrency configuration to latency-optimal TTS.
Speculative decoding
Speculative decoding uses a small draft model to generate several draft tokens autoregressively, and the target model to verify them in parallel while ensuring a lossless acceleration Leviathan et al. (2023); Chen et al. (2023). Many studies have proposed improvements to speculative decoding. EAGLE-series Li et al. (2024b, a, 2025), Medusa Cai et al. (2024), and Hydra Ankner et al. (2024) employ features from the target model for better draft acceptance rate. HASS Zhang et al. (2025) mitigates the inconsistency between training and inference by simulating the multi-step draft generation in the training phase. Our work demonstrates the feasibility of latency-aware TTS optimization through speculative decoding, presenting it as one viable approach to latency-optimal TTS.
3 Rethinking Test-Time Scaling with Latency
Previous works measure the budget of LLM test-time scaling by #tokens. However, it provides an incomplete evaluation, particularly when considering real-world deployment constraints. Our investigation reveals the necessity to consider latency-aware TTS in latency-sensitive scenarios.
Figure 4: Overview of how to improve TTS with latency budget. (a): default sequential scaling suffers from long latency due to low GPU utilization. (b): Branch-wise parallelism employs more branches to utilize FLOPs, improving accuracy with no extra time. (c): Sequence-wise parallelism utilizes FLOPs with speculative decoding, reaching faster generation and theoretically no performance loss. (d): With speculative decoding, the curve (red) shifts to the left (yellow), indicating reduced latency. With more branches, the curve shifts upward (blue), indicating improved performance. Latency-optimal TTS can be achieved by jointly applying these two optimizations (green).
3.1 Latency-Aware Test-time Scaling
The inference of LLMs on modern accelerators is a memory-bound process, constrained by memory bandwidth. Nowadays, LLMs have billions of parameters Bai et al. (2023); Yang et al. (2024); Team (2025); Guo et al. (2025); Meta (2024). In the autoregressive decoding, parameters (weights & KV caches) are loaded from memory (e.g., HBM on a GPU) into the compute units (e.g., SMs on a GPU) for matrix multiplication and other arithmetic operations. The runtime paid for the iteration is dominated by memory access. This is called the memory-bound nature of LLM. Under memory-bound constraints, slightly increasing the computation overhead does not affect the inference latency, but can increase throughput, as shown in Figure 3.
To characterize latency-aware test-time scaling under the memory-bound scenario, we explicitly consider two key factors:
Token efficiency is the ratio of task-specific accuracy improvement per token generated by the LLM, measured by %/token. A high token efficiency indicates that each token contributes significantly to achieving the accuracy. Prior work’s approach to determining compute-optimal TTS using #token as budget fundamentally represents a search for maximal token efficiency.
Throughput is the number of output tokens generated per wall-clock time, measured by token/s. Note that the throughput may vary as the sequence length grows. A high throughput indicates efficient use of hardware resources and the system’s capacity to quickly handle a large volume of tokens.
Consider s1.1-32B Muennighoff et al. (2025) as an example. Figure 1 compares the accuracy-#token and accuracy-latency curves of sequential and parallel scaling. Sequential scaling achieves higher token efficiency (83.3%/$2^{11.3}$toks) than parallel scaling (83.3%/$2^{14.4}$toks), but Figure 3 reveals its throughput is 16×16$\times$16 × lower. Consequently, sequential scaling requires 1.6×1.6$\times$1.6 × more time to attain comparable accuracy. This suggests that while sequential scaling improves token efficiency, its inferior throughput hinders better accuracy in limited time. Thus, test-time scaling budgets must consider both token efficiency and throughput to achieve high accuracy with low latency.
3.2 How to Improve Latency-Aware TTS? A Concurrency Perspective
To improve latency-aware TTS, the key insight is to improve generation concurrency to increase throughput. To this end, there exist two approaches from the perspective of concurrency:
Branch-wise Parallelism.
One approach is employing multiple concurrent branches $B$ for the question, as shown in Figure 4 (b). For instance, when a 2048-token response fails to yield a correct answer, users can infer more branches to generate multiple responses of the same length and determine the final answer through majority voting. This approach harnesses more underutilized memory-bound computational resources, introducing almost no extra latency. Employing multiple branches to explore diverse reasoning paths brings further improvements of TTS performance.
Sequence-wise Parallelism
Another effective approach is speculative decoding Leviathan et al. (2023); Chen et al. (2023), as shown in Figure 4 (c). SD accelerates by verifying multiple draft tokens concurrently, leveraging underutilized memory-bound computational resources to mitigate the memory access burden in sequential generation with lossless performance. With SD, LLM can generate longer responses within a limited time, thereby enhancing TTS performance.
The combination of the two approaches is illustrated in Figure 4 (d). Branch-wise parallelism enhances performance without increasing latency, elevating the scaling curve. Sequence-wise parallelism reduces latency without compromising performance, causing the scaling curve to shift leftward. Their combined effect moves the scaling curve toward the upper-left quadrant.
3.3 Latency-Optimal Test-Time Scaling
The joint application of both approaches increases the overall concurrency and introduces additional computational overhead. When this overhead surpasses memory access overhead, the system transitions into a compute-bound state, leading to more latency. Besides, branch-wise parallelism exhibits diminishing improvement with increasing branches, while sequence-wise parallelism’s acceleration reaches a maximum threshold. Consequently, an optimal boundary curve represents the latency-optimal test-time scaling, as shown by the green curve in Figure 4 (d). This curve defines where neither latency nor accuracy can be further improved without concurrency trade-offs.
We aim to determine the latency-optimal TTS strategy by allocating concurrency resources with parallel branches $B$ and draft length $\gamma$. Let $\text{Target}(\theta,T,x)$ be the output distribution over problem $x$ produced by the LLM with test-time compute hyperparameter $\theta$ and time limitation $T$. The settings of branches and draft length are included by $\{B,\gamma\}\subsetneq\theta$. The latency-optimal test-time scaling is given by:
$\theta^{*}_{x,y^{*}(x)}(T)=$
$\operatorname*{arg\,max}_{\theta}\left(\mathbb{E}_{y\sim\text{Target}(\theta,T,x)}\left[\mathbbm{1}_{y=y^{*}(x)}\right]\right),$ (1)
where $y^{*}(x)$ indicates the groundtruth of corresponding problem $x$, and $\theta^{*}_{x,y^{*}(x)}(T)$ represents the test-time latency-optimal scaling strategy for problem $x$ with time $T$. Finding the optimal $\theta^{*}_{x,y^{*}(x)}(T)$ is also the way to find optimal branches $B^{*}$ and draft length $\gamma^{*}$.
Figure 5: Latency-aware test-time scaling with different branches on s1.1-32B, LLama-3.1-8B-Instruct and s1.1-3B. Speculative decoding is not implemented. Each picture shows the scaling curve with varying branch numbers.
Figure 6: Latency-aware test-time scaling with speculative decoding of different draft length $\gamma$ on MATH-500. Number of branch is fixed to 1. Draft model: DeepSeek-R1-Distill-Qwen-7B. Target model: QwQ-32B.
Figure 7: Latency-aware test-time scaling curves on MATH-500 with sequential scaling in red, parallel scaling in blue, and latency-optimal in green (the same to subsequent figures).
Figure 8: Latency-aware test-time scaling curves of QwQ-32B and DeepSeek-R1-Distill-Qwen-32B under latency on AIME24 (left two panels) and AIME25 (right two panels).
4 Experiments and Discussion
4.1 Experimental Setup
Tasks.
Representative challenging tasks as benchmarks to measure the scaling property are selected: MATH-500 Hendrycks et al. (2021) is a popular math benchmark comprising 500 high-school competition problems. AIME24 AoPS (2024) and AIME25 AoPS (2025) each consist of 30 math problems from the 2024 and 2025 American Invitational Mathematics Examination (AIME). GPQA-Diamond Rein et al. (2024) consists of 198 science QA problems encompassing PhD-level physics, chemistry, and biology.
Models.
We choose models with parameter sizes suitable for device-side deployment. Also, to implement speculative decoding, we select model types with draft models available. Therefore, considering its scalability, we employ s1.1-32B as our test-time scaling baseline. s1.1-7B is used as its draft model for speculative decoding. Also, we conduct relevant experiments on LLaMa-3.1-8B-Instruct Meta (2024) with Eagle3 Li et al. (2025) as the draft model. For RL-based thinking model, we employ DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025) and QwQ-32B Team (2025) for their superior reasoning ability. The DeepSeek-R1-Distill-Qwen-7B is used as their draft model. We employ s1.1-3B for small-sized LLM for its outstanding model size and excellent performance with Qwen2.5-0.5B-Instruct Bai et al. (2023) as the draft model. To better control the output sequence length, we employ budget forcing Muennighoff et al. (2025) as an appropriate method to enforce the model to generate longer CoT. We use majority voting Wang et al. (2022) to select the answer. Experiments are conducted on the codebase of OpenR Wang et al. (2024a).
4.2 Sequential and Parallel Scaling under Latency-Aware TTS
Sequential scaling suffer from low throughput.
As shown in Figure 2, sequential scaling achieves superior performance across most model types. However, its impact is less pronounced for models not explicitly trained for long CoT reasoning, such as LLaMa-3.1-8B-Instruct Meta (2024). Conversely, Figure 3 reveals that sequential scaling exhibits lower throughput. For instance, the s1.1-32B model processes only 22.7 tokens/s, translating to approximately 1,000 tokens per minute. While sequential scaling demonstrates better token efficiency, its lower throughput makes it less effective than parallel scaling under time constraints. Consequently, sequential scaling struggles to generate sufficient tokens to attain high accuracy.
Parallel scaling improves fast, but limited.
In Figure 2, parallel scaling shows lower token efficiency than sequential scaling for most models except LLaMa-3.1-8B-Instruct. However, from Figure 3, increasing the branches of LLM inference hardly requires extra latency. Specifically, when increasing branches from 1 to 16, s1.1-32B gains nearly $10\%$ improvement on MATH-500 at almost no extra latency cost. This can be attributed to the fact that the computation resources are fully utilized. By inference in branches, one autoregressive decoding procedure would generate the branches’ tokens, but the latency cost of developing one token by sequential scaling is the same. So parallel scaling can reach a larger #token faster. When the branches are increased to 64, the performance gain becomes slight, and the latency slightly grows because of the large amount of KV cache. Overall, parallel scaling maximizes hardware parallelism and scales output tokens simultaneously. But, parallel scaling often yields suboptimal accuracy due to branch redundancy and limited scalability.
Mostly, parallel scaling can surpass sequential scaling within limited time period.
As shown in Figure 2, merely employing parallel scaling can surpass sequential scaling within a shorter, limited time. However, the extent of performance improvement varies depending on the model type. Specifically, for reasoning models like QwQ-32B, parallel scaling reaches $80.1\%$ accuracy on the MATH-500 dataset within $30$ seconds, but sequential scaling requires 1.4×1.4$\times$1.4 × more time to achieve the comparable performance. Sequential scaling shows a slight performance gain for LLaMa-3.1-8B-Instruct, which is not designed for long CoT reasoning. While parallel scaling is still effective on MATH-500 since it explores more diverse solution paths, it shows obvious token efficiency, and sequential scaling cannot surpass it. Overall, parallel scaling can surpass sequential scaling w.r.t. latency, reaching a comparable accuracy within a relatively short time.
4.3 The Impact of Branch-wise Parallelism for Latency-Aware TTS
We conduct experiments on sequential scaling with varying branch sizes. In these configurations, we implement sequential scaling in parallel, with all parallel branches aggregated through majority voting. As shown in Figure 5, s1.1-32B demonstrates an initial upward trend in the TTS curve as the number of branches increases, indicating effective performance improvements. However, when the branch size grows excessively, the performance gains diminish, and latency increases slightly due to the non-negligible overhead of KV cache. In contrast, LLaMa-3.1-8B-Instruct exhibits minimal sequential scaling effects, resulting in a flatter curve compared to s1.1-32B and QwQ. Nevertheless, parallel scaling proves impactful, yielding an accuracy improvement that elevates the curve.
Strategy 1024 2048 4096 8192
Lat. (s) Acc. (%) Lat. (s) Acc. (%) Lat. (s) Acc. (%) Lat. (s) Acc. (%)
Baseline 96.2 $\pm$0.2 75.0 $\pm$0.6 157.9 $\pm$0.5 83.4 $\pm$1.2 280.6 $\pm$0.5 84.9 $\pm$1.8 419.5 $\pm$0.7 85.4 $\pm$0.9
Bnh-wise 96.7
$\pm$0.2
79.4
$\pm$0.7
159.2
$\pm$0.8
86.2
$\pm$1.1
284.5
$\pm$0.3
88.3
$\pm$1.5
428.0
$\pm$0.6
89.6
$\pm$1.0
Seq-wise 55.5
$\pm$0.4
75.0
$\pm$0.3
91.1
$\pm$0.3
83.4
$\pm$1.9
161.9
$\pm$0.3
84.9
$\pm$1.4
241.9
$\pm$0.8
85.4
$\pm$0.1
Lat-Opt. 57.2 $\pm$0.4 82.3 $\pm$1.8 95.9 $\pm$0.3 85.9 $\pm$0.7 179.1 $\pm$0.4 91.7 $\pm$1.3 283.1 $\pm$0.6 92.7 $\pm$0.3
Improve 1.7$\times$ 7.3+ 1.6$\times$ 2.5+ 1.6$\times$ 6.8+ 1.5$\times$ 7.3+
Table 1: Results of baseline, branch-wise parallelism, sequence-wise parallelism and latency-optimal TTS with different sequence lengths on MATH-500 of s1.1-32B. Lat.: Latency. Acc.: Accuracy. Bnh-wise: Branch-wise parallelism. Seq-wise: Sequence-wise parallelism. Lat-Opt.: Latency-Optimal. Improve is reported between Baseline and Lat-Opt. Results are obtained from 3 repeated experiments with mean and standard deviation reported.
Figure 9: Latency-aware TTS curves of s1.1-32B on MATH-500, AIME24, AIME25 and GPQA-Diamond.
Figure 10: Latency-aware TTS curves of s1.1-3B on MATH-500, AIME24, AIME25 and GPQA-Diamond.
4.4 The Impact of Sequence-wise Parallelism for Latency-Aware TTS
Speculative decoding can accelerate model inference to some extent while preserving accuracy. As shown in Figure 6, most scaling curves with speculative decoding are on the left side of the baseline curve. As the speed-up ratio grows, the left-shifted trend becomes more obvious. However, when the draft length $\gamma$ becomes large, the speed-up ratio would decay, slowing down the overall speed. In our experiments, QwQ-32B is the target model, while DeepSeek-R1-Distill-Qwen-7B as the draft model, achieving a maximum speed-up ratio of 1.64×1.64$\times$1.64 ×. Therefore, the scaling curve moves to the left first and then to the right. Sequential scaling with speculative decoding can achieve $7.5\%$ higher accuracy than the baseline curve for a limited time. These results demonstrate that employing a proper draft length $\gamma$ can push the inference speed to an optimal stage, enabling further sequential scaling within a limited time.
4.5 Latency-Optimal TTS with Branch-wise and Sequence-wise Parallelism
We conduct comprehensive experiments with various parallel branches and draft lengths to identify the concurrency configuration to achieve latency-optimal TTS. The results are shown in Figure 7 and 8. For LLMs that benefit primarily from sequential scaling, speculative decoding emerges as the dominant factor in the latency-optimal configuration. Conversely, for LLMs hat exhibit improvements from branches, increasing the number of branches yields huge performance gains. This difference stems from the different token efficiency. For reasoning models like QwQ, accuracy improvements are achieved through long CoT, making better SD acceleration more advantageous within limited time. In contrast, models like LLaMA-3.1-8B-Instruct, which do not benefit from long CoT, using more branches to expand the search paths is more efficient. See appendix for detailed configurations.
5 Results under Latency-Optimal TTS
5.1 Can LLM Solve the Problem in 1 minute?
Achieving as high accuracy as possible within a limited time holds significant practical value. We aim to find out how TTS can improve the accuracy individually on the device side within a relatively short time limitation, like 1 minute. To this end, we comprehensively evaluate s1.1-32B and s1.1-3B w.r.t. latency and record their inference latency.
From Figures 10 and 10, we derive the following observations: (1) Large model like s1.1-32B, achieves relatively high accuracy within just 1 minute, which is unattainable by the baseline. (2) Smaller model (s1.1-3B) attains notable accuracy in merely 10 seconds, showing significant potential. This trend persists across datasets spanning diverse domains, suggesting that popular models can be optimized using the latency-optimal TTS strategy.
5.2 How Can Latency-Optimal Improve Compared with Baseline?
Based on previous findings of latency-optimal TTS with different model types, branches, and draft lengths, we summarize the results in Table 1. For s1.1-32B on MATH-500, we find that latency-optimal TTS can achieve $6\%$ accuracy improvement on average than merely using sequence-wise parallelism, and 1.6×1.6$\times$1.6 × on average faster than branch-wise parallelism. Latency-optimal TTS outperforms the baseline in both accuracy and latency to a noticeable degree. However, increased branch and draft length causes LLMs to enter a compute-bound regime, where computational overhead exceeds memory access overhead. This eliminates the benefits of parallelism, leading to increased latency and sometimes even worse than the baseline. This suggests that a latency-optimal TTS strategy requires extremely fine-grained parameter tuning.
6 Conclusion
In this paper, we propose to rethink test-time scaling in latency-sensitive scenarios. We show that compute-optimal does not always result in latency-optimal under such conditions due to memory-bound constraint. To address this, we propose to improve TTS on generation concurrency to maximize throughput from a unified view. Specifically, we present two approaches: (1) branch-wise parallelism via multiple branches and (2) sequence-wise parallelism by speculative decoding, along with their combinations. Furthermore, we investigate the concurrency allocation strategy to balance these approaches for latency-optimal TTS. Experimental results show that latency-optimal TTS enables 32B model to achieve 82.3% accuracy on MATH-500 within 1 minute, while 3B model attains notable accuracy within just 10 seconds, showing significant improvements in both speed and accuracy.
7 Limitations
The basic concept of this paper is that the inference of LLM is a memory-bound process. However, this concept holds on small (like mobile phones, personal computers) or medium (like work stations) scale hardware. For large-scale servers, which deals with hundreds or thousands of requests at one time, the main bottleneck of inference is computation, and the budget of inference can be measured by #tokens. However, the studies on small and medium scale hardware still hold significant meaning, as the practical budget measured on these platforms is often under memory-bound scenarios.
The maximum sequence length in our experiment settings is 8192. Note that, as the number of branches and sequence length increase, the substantial memory access burden imposed by the KV cache leads to a significant rise in latency. This causes methods that increase concurrency to experience corresponding increases in latency. Additionally, longer sequence lengths may introduce performance degradation due to excessive context length in the model Muennighoff et al. (2025). Since our work adopts relatively common experimental configurations, these issues have not manifested prominently, and thus our proposed method and conclusions remain valid.
Appendix A Appendix
A.1 Why does the parallel scaling curve exhibit a steeper slope than the sequential scaling?
This result can be attributed to the heavy model weights of LLM under the measurement of memory access. For instance, Qwen2.5-32B-Instruct has 32 billion parameters, so the model weights is 64GB. While the total KV cache at sequence length 1024 and batch size 1 is 0.25GB. Under this condition, KV cache only accounts for a small portion of memory access. When parallel scaling can easily extend branches with a small budget by 0.25GB per branch with 1024 length, sequential scaling tries hard to load the whole model weights of the entire 64GB just for 1 more extended token. This extreme imbalance on memory access makes parallel scaling significantly advantageous regarding memory access. In contrast, sequential scaling suffers from high memory access costs.
A.2 Additional Results
We present additional experimental results. The results on the influence of branch-wise parallelism is shown in Figure 11. The results of the influence of branch-wise parallelism under different speculative decoding configurations is shown in Figure 12. The results of sequence-wise parallelism under different branch counts are shown in Figure 13. The results of the combined branch-wise parallelism and sequence-wise parallelism configurations w.r.t. fixed generation concurrency $F$ are shown in Figure 14.
The acceptance rate $\alpha$ of the speculative decoding we employ is reported below: Target: s1.1-32B, draft: s1.1-7B, $\alpha$: 0.831. Target: DeepSeek-R1-Distill-Qwen-32B, draft: DeepSeek-R1-Distill-Qwen-7B, $\alpha$: 0.897. Target: QwQ-32B, draft: DeepSeek-R1-Distill-Qwen-7B, $\alpha$: 0.781. Target: LLaMa-3.1-8B-Instruct, draft: Eagle3, $\alpha$: 0.904. Target: s1.1-3B, draft: Qwen2.5-0.5B-Instruct, $\alpha$: 0.701.
The latency-optimal configurations shown in Figure 7,  8,  10 and 10 are reported below: Figure 7: Left: B=16, $\gamma$=5. Right: B=32, $\gamma$=3. Right: B=32, $\gamma$=5. Figure 8: AIME24 with QwQ-32B: B=16, $\gamma$=4. AIME24 with DeepSeek-R1-Distill-Qwen-32B: B=32, $\gamma$=4. AIME25 with QwQ-32B: B=16, $\gamma$=4. AIME24 with DeepSeek-R1-Distill-Qwen-32B: B=32, $\gamma$=5. Figure 10: MATH-500: B=16, $\gamma$=5. GPQA-Diamond: B=32, $\gamma$=5. AIME24: B=32, $\gamma$=4. AIME25: B=32, $\gamma$=5. Figure 10: MATH-500: B=32, $\gamma$=5. GPQA-Diamond: B=16, $\gamma$=5. AIME24: B=8, $\gamma$=5. AIME25: B=16, $\gamma$=5.
Figure 11: More results of the influence of the number of branches from branch-wise parallelism.
Figure 12: More results of the influence of the number of branches from branch-wise parallelism under different sequence-wise configurations. The draft lengths of speculative decoding varies among {2, 4, 5}.
Figure 13: More results of the influence of draft length from sequence-wise parallelism under different branch-wise configurations. The number of branches varies among {1, 4, 16, 64}.
Figure 14: Results of different branch-wise parallelism and sequence-wise parallelism configurations with fixed generation concurrency.